OpenStudy (anonymous):

Dimensional Analysis Question convert 3.42g into kilograms

OpenStudy (anonymous):

In order to conver grams to kilograms you must divide by 1000. There 3.42g/1000=.00342

OpenStudy (australopithecus):

\[\frac{1kg}{1000g}3.42g = \frac{(1kg)3.42g}{1000g} = \frac{1*3.42(g)kg}{1000g} = \frac{3.42kg}{1000}\]

OpenStudy (australopithecus):

1kg = 1000g

OpenStudy (australopithecus):

if you wanted to convert back to grams you would just flip the ratio so that kg cancels out